i fogot the steps to eliminations.
0.3x - 0.2y = 4
0.5x + 0.3y = - 7/17
can someone help me

Answer by ilana(307)

To use the elimination method, make either the x terms or the y terms cancel out when you add or subtract the equations.
You can start by multiplying both equations by 10 to get rid of the decimals.
3x - 2y = 40
5x + 3y = -70/17
Multiply the first equation by 3 and the second equation by 2 so you can add -6y and 6y to get 0.
9x - 6y = 120
10x + 6y = -140/17
[I got -140/17 by multiplying -70/17 by 2. Anything you do to one side of the equation you have to do to the other in order for the equation to still work. So we multiplied each term on each side by 2.]
Now add the equations.
19x = 120 - 140/17
19x = (2040 - 140)/17
19x = 1900/17
x = 1900/(17*19)
x = 1900/323
Now plug this x into either equation to solve for y.
3x - 2y = 40
3(1900/323) - 2y = 40
5700/323 - 2y = 40
2y = 5700/323 - 40
2y = (5700 - 12920)/323
2y = -7220/323
y = -7220/646
The approximate solution is (5.88, -11.18).